# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

### v1

@@ Appendix B Proofs of Lemmas 2, 5, and 8
Define $\mathsf{W}_{i}=\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)$ and $\tilde{\mathsf{W}}_{i}=\mathsf{W}_{i}-\mathbb{E}\mathsf{W}_{i}$. Then for the decoding metric (6) of CDF-ORBGRAND, it follows that
@@
For $i\neq j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]-\mathbb{E}[\mathsf{W}_{i}]\mathbb{E}[\mathsf{W}_{j}]$. We now turn to the evaluation of $\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$, which, by definition, can be expressed as (37), shown at the top of this page. The first term in (37) leads to (38) and (39), shown at the top of this page and the next page, respectively. Note that $\mathsf{R}_{i}$ and $\mathsf{R}_{j}$ in (39) are the ranks resulting from inserting $|t_{i}|=\Bigl\lvert\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}\Bigr\rvert$ and $|t_{j}|=\Bigl\lvert\ln\frac{q^{+}(y_{j})}{q^{-}(y_{j})}\Bigr\rvert$ into a sorted array of N2N-2 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y_{i},y_{j})$ the upper branch of (39). Therefore, (38) can be rewritten as $\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$. Applying a similar procedure to the other three terms in (37), we obtain
@@
$\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$ $=\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)\mathbf{1}(\text{sgn}(\mathsf{T}_{j})\cdot\mathsf{X}_{j}(1)<0)\Big]$
@@
$=\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\Big]$
@@
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=+1\Big]$
@@
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}>0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=-1\Big]$
@@
$\quad+\frac{1}{4}\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}>0)\Big\lvert\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=-1\Big],$ (37)
@@
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\Big]$
@@
$=\quad$ $\mathbb{E}\Big[\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i},\mathsf{Y}_{j}\Big]\Big],$ (38)
@@
$\mathbb{E}\Big[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Big\lvert\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i}=y_{i},\mathsf{Y}_{j}=y_{j}\Big]$
@@ C-A Proof of Lemma 3
$\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\mathbb{E}\left\{\prod_{i=1}^{N}e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\prod_{i=1}^{N}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\},$ (41)
where we use the fact that $\underline{\mathsf{T}}$ is induced by $\underline{\mathsf{X}}(1)$ and is therefore independent of $\underline{\mathsf{X}}(w^{\prime})$. For each product term in (41), we have
$\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Big\lvert\underline{\mathsf{T}}\right\}$
$=\quad$ $\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}<0)}\Big\lvert\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=+1\right\}$
$+\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}>0)}\Big\lvert\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=-1\right\}$
@@
$=\quad$ $\frac{1}{2}\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\Bigr),$ (42)
where since $\mathsf{R}_{i}$ is determined by $\underline{\mathsf{T}}$, the conditional expectation can be removed. Therefore, we have
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\Bigr)-\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\Bigr)-\ln 2,$ (43)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\lim\limits_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\ln\Bigl(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\Bigr)-\ln 2$
$=\quad$ $\int_{0}^{1}\ln\Bigl(1+e^{\theta\Psi^{-1}(u)}\Bigr)\mathrm{d}u-\ln 2$
$=\quad$ $\int_{0}^{\infty}\ln\Bigl(1+e^{\theta t}\Bigr)\Psi^{\prime}(t)\mathrm{d}t-\ln 2$
$=\quad$ $\mathbb{E}\Bigl[\ln\Bigl(1+e^{\theta\cdot\Bigl\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-\ln 2,$ (44)
@@ C-B Proof of Lemma 6
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{m}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{\mathsf{R}_{i,j}}{mN+1}\right)}\Bigr)-m\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{mN}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{n}{mN+1}\right)}\Bigr)-m\ln 2.$ (45)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big\lvert\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $m\int_{0}^{1}\ln\Bigl(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}(u)}\Bigr)\mathrm{d}u-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\bar{\Psi}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\cdot\Bigl(\frac{1}{m}\sum_{j=1}^{m}\Psi_{j}^{\prime}(t)\Bigr)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\int_{0}^{\infty}\ln\Bigl(1+e^{\frac{\theta}{m}t}\Bigr)\Psi_{j}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\mathbb{E}\Bigl[\ln\Bigl(1+e^{\frac{\theta}{m}\cdot\Bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\Bigr\rvert}\Bigr)\Bigr]-m\ln 2,$ (46)

### v2

@@ Appendix B Proofs of Lemmas 2, 5, and 8
Define $\mathsf{W}_{i}=\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)$ and $\tilde{\mathsf{W}}_{i}=\mathsf{W}_{i}-\mathbb{E}[\mathsf{W}_{i}]$. Then for the decoding metric (6) of CDF-ORBGRAND, it follows that
@@
For $i\neq j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]-\mathbb{E}[\mathsf{W}_{i}]\mathbb{E}[\mathsf{W}_{j}]$. We now turn to the evaluation of $\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$, which, by definition, can be expressed as (37), shown at the top of this page. The first term in (37) leads to (38) and (39), shown at the top of this page and the next page, respectively. Note that $\mathsf{R}_{i}$ and $\mathsf{R}_{j}$ in (39) are the ranks resulting from inserting $\lvert t_{i}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}\bigr\rvert$ and $\lvert t_{j}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{j})}{q^{-}(y_{j})}\bigr\rvert$ into a sorted array of N2N-2 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y_{i},y_{j})$ the upper branch of (39). Therefore, (38) can be rewritten as $\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$. Applying a similar procedure to the other three terms in (37), we obtain
@@
$\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$ $=\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(1)<0)\mathbf{1}(\text{sgn}(\mathsf{T}_{j})\cdot\mathsf{X}_{j}(1)<0)\right]$
@@
$=\frac{1}{4}\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\right]$
@@
$\quad+\frac{1}{4}\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}<0)\Bigg|\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=+1\right]$
@@
$\quad+\frac{1}{4}\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}>0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=-1\right]$
@@
$\quad+\frac{1}{4}\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}>0)\mathbf{1}(\mathsf{T}_{j}>0)\Bigg|\mathsf{X}_{i}(1)=-1,\mathsf{X}_{j}(1)=-1\right],$ (37)
@@
$\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1\right]$
@@
$=\quad$ $\mathbb{E}\left[\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i},\mathsf{Y}_{j}\right]\right],$ (38)
@@
$\mathbb{E}\left[\Psi^{-1}\Big(\frac{\mathsf{R}_{i}}{N+1}\Big)\Psi^{-1}\Big(\frac{\mathsf{R}_{j}}{N+1}\Big)\cdot\mathbf{1}(\mathsf{T}_{i}<0)\mathbf{1}(\mathsf{T}_{j}<0)\Bigg|\mathsf{X}_{i}(1)=+1,\mathsf{X}_{j}(1)=+1,\mathsf{Y}_{i}=y_{i},\mathsf{Y}_{j}=y_{j}\right]$
@@ C-A Proof of Lemma 3
$\mathbb{E}\left\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\right\}$
$=\quad$ $\mathbb{E}\left\{\prod_{i=1}^{N}e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Bigg|\underline{\mathsf{T}}\right\}$
$=\quad$ $\prod_{i=1}^{N}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Bigg|\underline{\mathsf{T}}\right\},$ (41)
where we use the fact that $\underline{\mathsf{T}}$ is induced by $\underline{\mathsf{X}}(1)$ and is therefore independent of $\underline{\mathsf{X}}(w^{\prime})$. Each term in (41) is
$\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\text{sgn}(\mathsf{T}_{i})\cdot\mathsf{X}_{i}(w^{\prime})<0)}\Bigg|\underline{\mathsf{T}}\right\}$
$=\quad$ $\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}<0)}\Bigg|\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=+1\right\}$
$+\frac{1}{2}\mathbb{E}\left\{e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)\cdot\mathbf{1}(\mathsf{T}_{i}>0)}\Bigg|\underline{\mathsf{T}},\mathsf{X}_{i}(w^{\prime})=-1\right\}$
@@
$=\quad$ $\frac{1}{2}\left(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\right),$ (42)
where since $\mathsf{R}_{i}$ is determined by $\underline{\mathsf{T}}$, the conditional expectation can be suppressed. Therefore, we have
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\ln\left(1+e^{\theta\Psi^{-1}\left(\frac{\mathsf{R}_{i}}{N+1}\right)}\right)-\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{N}\ln\left(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\right)-\ln 2,$ (43)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\lim\limits_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\ln\left(1+e^{\theta\Psi^{-1}\left(\frac{n}{N+1}\right)}\right)-\ln 2$
$=\quad$ $\int_{0}^{1}\ln\left(1+e^{\theta\Psi^{-1}(u)}\right)\mathrm{d}u-\ln 2$
$=\quad$ $\int_{0}^{\infty}\ln\left(1+e^{\theta t}\right)\Psi^{\prime}(t)\mathrm{d}t-\ln 2$
$=\quad$ $\mathbb{E}\left[\ln\left(1+e^{\theta\cdot\bigl\lvert\ln\frac{q^{+}(\mathsf{Y})}{q^{-}(\mathsf{Y})}\bigr\rvert}\right)\right]-\ln 2,$ (44)
@@ C-B Proof of Lemma 6
$\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{m}\ln\left(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{\mathsf{R}_{i,j}}{mN+1}\right)}\right)-m\ln 2$
$=\quad$ $\frac{1}{N}\sum_{n=1}^{mN}\ln\left(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}\left(\frac{n}{mN+1}\right)}\right)-m\ln 2.$ (45)
@@
$\lim\limits_{N\to\infty}\frac{1}{N}\ln\mathbb{E}\Bigl\{e^{N\theta\mathsf{D}(w^{\prime})}\Big|\underline{\mathsf{T}}\Bigr\}$
$=\quad$ $m\int_{0}^{1}\ln\left(1+e^{\frac{\theta}{m}\bar{\Psi}^{-1}(u)}\right)\mathrm{d}u-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\left(1+e^{\frac{\theta}{m}t}\right)\bar{\Psi}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $m\int_{0}^{\infty}\ln\left(1+e^{\frac{\theta}{m}t}\right)\cdot\Bigl(\frac{1}{m}\sum_{j=1}^{m}\Psi_{j}^{\prime}(t)\Bigr)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\int_{0}^{\infty}\ln\left(1+e^{\frac{\theta}{m}t}\right)\Psi_{j}^{\prime}(t)\mathrm{d}t-m\ln 2$
$=\quad$ $\sum_{j=1}^{m}\mathbb{E}\Biggl[\ln\Biggl(1+e^{\frac{\theta}{m}\cdot\bigl\lvert\ln\frac{q_{j}^{+}(\mathsf{Y})}{q_{j}^{-}(\mathsf{Y})}\bigr\rvert}\Biggr)\Biggr]-m\ln 2,$ (46)